# i-PI 3.0: a flexible, efficient framework for advanced atomistic simulations

### v1

@@ Abstract
Atomic-scale simulations have progressed tremendously over the past decade, largely due to the availability of interatomic potentials based on machine-learning algorithms. These potentials enable the combination of the accuracy of electronic structure calculations with extensive length and time scales. In this paper, we present a new release of the i-PI code that allows the community to fully benefit from the rapid developments in the field. i-PI is a Python software that facilitates the integration of different methods and different software tools by using a socket interface for inter-process communication. The current framework is flexible enough to support rapid prototyping and the combination of various simulation techniques, while maintaining a speed that prevents it from becoming the bottleneck in most workflows. We discuss the implementation of several new features, including an efficient algorithm to model bosonic and fermionic exchange, a framework for uncertainty quantification to be used in conjunction with machine-learning potentials, a communication infrastructure that allows deeper integration with electronic-driven simulations, and an approach to simulate coupled photon-nuclear dynamics in optical or plasmonic cavities. For this release, we also improved some computational bottlenecks of the implementation, reducing the overhead associated with using i-PI over a native implementation of molecular dynamics techniques. We show numerical benchmarks using widely adopted machine learning potentials, such as Behler-Parinello, DeepMD and MACE neural networks, and demonstrate that such overhead is negligible for systems containing between 100 and 12000 atoms.
@@ I Introduction
The atomic-scale modelling of molecules and materials is a constantly evolving field driven by the interest of researchers in discovering new fundamental phenomena, interpreting complex experimental results, and, together with an increasing number of commercial enterprises, improving, discovering, and developing new materials.[1, 2, 3, 4, 5, 6] Several computational developments in recent years facilitated an extraordinary advancement of the problems that can be addressed by atomistic simulations, far exceeding the steady improvement of the underlying hardware. Arguably, the largest contribution to such fast progress can be attributed to the emergence of machine-learning interatomic potentials (MLIPs). MLIPs have cut down the computational cost of calculations with ab initio accuracy and allowed access to system sizes and time scales which were previously only accessible to empirical interatomic potentials.[7, 8, 9, 10] This has driven a renewed interest in complex simulation algorithms that were thought inapplicable for systems that require an ab initio quality description of their interatomic interactions, such as the study of catalytic reactions [11, 12, 13] or thermodynamic stability of materials.[14, 15, 16] However, the lack of easy-to-use implementations of such approaches has become an important factor that hinders their adoption by many researchers’ workflows.
@@
Since its first release almost 10 years ago, one of i-PI’s primary objectives has been “lowering the implementation barrier to bring state-of-the-art sampling and atomistic modelling techniques to ab initio and empirical potentials programs”.[17] The first iteration of i-PI was tailored to perform path integral molecular dynamics (PIMD) simulations, hence the code name. Since then, and thanks to a community effort, it has evolved into a “universal force engine for advanced molecular simulations” including standard and advanced simulations algorithms, as described in the following sections.[18]
@@
In the last few years, our efforts have adapted to the changing needs of the community. While continuing with the implementation of new and advanced features, we have improved the scalability of the code backend, as required by newly available and computationally efficient MLIPs. We also adopted automatic testing workflows to strengthen the code stability and enhance our reproducibility standards.
@@
i-PI’s capabilities as a versatile tool for atomistic simulations can be demonstrated by the wide range of applications in which it has been employed. These examples range from the dynamics of charge density wave materials[19] and nuclear motion associated to electronic excitations [20], to the elucidation of the thermodynamic properties of aqueous electrons,[21] the calculation of first principles phase diagrams,[22, 23] and the study of aqueous solutions and interfaces.[24, 25, 26, 27] Moreover, by virtue of its modular structure, i-PI can be used as a testbed to combine different simulation “modes” to design specialized protocols tailored to the user’s requirements. We are convinced that i-PI has the potential to empower many more researchers in chemistry, physics and materials science to adopt cutting-edge simulation techniques in their research.
@@
This manuscript reports on the progress made in the last 5 years since the release of version 2.0, by describing the newly adopted maintenance strategies, the improved performance, and the newly implemented features. The manuscript is organised as follows: In Sec. II we start with a short review of the code architecture followed by a description of the improvements made to the code’s robustness and usability and a list of main features available in this release. In Sec. III, we present a detailed discussion of the code efficiency for production and large-scale simulations, while in Sec. IV, we describe the new features since the last code release along with some numerical examples. In Sec. V, we showcase how the modularity and flexibility of i-PI help the implementation of advanced simulation setups that would otherwise require new codes to be written and Sec. VI concludes the article.
@@ II Program Overview
i-PI  is a force engine that operates with a client-server paradigm. i-PI  acts as a server in command of the evolution of the nuclear positions, while one or more instances of the client code handle the evaluation of system-specific properties. Designed to be universal, i-PI’s architecture is agnostic to the indentity of the force providers, with a general and flexible backend that accommodates to a wide range of client codes.
@@
The code is written in Python 3, a high-level, general-purpose interpreted programming language known for its emphasis on code readability. This choice facilitates rapid prototyping of new ideas and relatively easy code maintenance when compared to compiled languages traditionally used in scientific computing. Additionally, the availability of highly optimised numerical libraries allows Python-based software to deliver competitive performance. Indeed, even computing-intensive electronic structure codes have released Python implementations in recent years.[28, 29] A detailed discussion of the code efficiency, and improvements since the previous release, is presented in Sec. III.
@@
The communication between i-PI  and the client codes is carried out through sockets, enabling local connections by shared-memory inter-process communication (UNIX domain sockets) or between two hosts connected to a local network or the Internet (TCP/IP sockets). During the initialization, i-PI  creates the sockets and awaits the connection of one or more clients. When the connection is established, the simulation loop starts: i) i-PI communicates the nuclear positions and the cell lattice vector to the active clients, ii) the clients evaluate the required properties by the chosen simulation algorithm, such as energy, forces, stresses, etc, and send these back, and iii) i-PI uses this data to displace the atoms and to calculate the new nuclear positions, restarting the loop. As i-PI  is designed to be client-agnostic, clients must be initialized independently before they can connect to the sockets. Importantly, the identity and number of the atoms remain constant throughout the simulations, so the geometries used to initialize i-PI  and the client code must be consistent.
@@
A schematic representation of the code structure and the server-client communication is presented in Fig. 1. i-PI is structured in a modular way that represents the underlying physics of the target simulation as faithfully as possible. To achieve this, the code is organized around the System class which encodes all the information related to the physical system, such as the number and identity of atoms, the ensemble to be sampled, the number of replicas in path integral simulations, the initialization procedure, and the algorithm for evolving the nuclear positions. Forces are managed through the Forces class, which integrates the individual force components, each of which is computed following the strategy specified in a Forcefield class. This two-layer approach is particularly advantageous for algorithms where the total forces and energies are obtained by a combination of these quantities computed at different accuracy levels,[30, 31, 32, 33, 34] or among different system portions.[35] Simulation techniques that require the evolution of many systems simultaneously make use of the SMotion (Systems Motion) class. This class enables the definition of evolution steps that combine the (possibly many) different systems, facilitating, for example, exchanges between system replicas as done in replica exchange simulations. Finally, estimators can be defined within the code or computed by the provided post-processing tools or user-generated scripts.
For the sake of completeness, we report an updated list of the code features in section II.4.
II.1 Why you might want to get involved: a developer perspective
There are numerous reasons for contributing to i-PI  that developers in the field might find compelling. We summarise some of the key arguments below.
Many available clients codes:
@@ II.1 Why you might want to get involved: a developer perspective
i-PI  is connected to a multitude of electronic-structure and machine learning codes, either directly or through its ASE [36] or LAMMPS [37] interfaces. This wide-ranging connectivity means that a single implementation is readily available to all the users of these codes, which significantly amplifies the potential impact of each contribution to i-PI .
@@
Similarly, in scenarios where a developer aims to use a newly implemented algorithm with different types of codes, such as a density functional theory (DFT) plane wave code to study periodic systems, or a quantum chemistry code to study gas phase molecules at a higher electronic structure level, two separate implementations would be required. By integrating the algorithm into i-PI, only one implementation is required and there is only one code to maintain.
@@
i-PI ’s infrastructure automatically manages updates of energy, forces, stresses, and other structure-dependent properties by establishing dependencies between physically related variables. This enables developers to concentrate solely on the algorithm’s implementation making the development process more focused and efficient.
@@ II.2 Code robustness and usability
The significant increase of i-PI  users and contributions over the last years, primarily from scientists outside the core developers’ research groups, required enhancing the maintainability, robustness, and standardization of the code. These improvements involve five distinct aspects:
@@
The documentation of the i-PI code is now generated with Sphinx and available online. It is easily accessible at https://ipi-code.org/i-pi/index.html and linked to the main code webpage. The documentation has been updated to reflect the current status of the code, and a section about how to contribute new features to the code has been added to reduce the barrier for new developers. An automatic parsing of the doc strings throughout i-PI ensures a complete listing of all possible keywords and the corresponding explanation, which is automatically updated as soon as new features are implemented and merged in the main branch.
@@
We have completely restructured the provided examples to make them as intuitive as possible for newcomers. The examples are now organized into three folders, namely features, clients, and hpc_script. The features folder displays extensive examples of the different implemented functionalities, all of which can be run locally using the built-in Python or Fortran driver, without requiring additional software installation. The clients folder includes examples tailored to specific driver codes, highlighting the unique syntax and tags needed for an adequate setup when used in connection to i-PI. At the moment, we provide examples for ASE,[36] CASTEP,[38] CP2K,[39] LAMMPS,[37] DFTB+,[40] ffsGDML,[41] FHI-aims,[42] librascal,[43] Quantum Espresso,[44, 45, 46] VASP,[47] elphmod,[48] and YAFF.[49] We note, however, that many other codes can also be connected to i-PI  through their ASE and LAMMPS interfaces. Finally, motivated by the number of user queries on this topic, the hpc_scripts folder provides examples with the corresponding submission scripts to run the simulations on high-performance computing platforms.
@@
We have added more complex examples to illustrate the application of i-PI  in realistic simulation scenarios, combining several of the features showcased in the examples folder. Some of these demonstration examples also cover some of the use of the pre- and post-processing scripts accompanying the code. Users specifically interested in learning the theory of path integral simulations and obtaining hands-on practice using i-PI  are referred to the online courses developed with this goal https://tinyurl.com/pimd-mooc or https://github.com/i-pi/piqm2023-tutorial/.
@@
We have implemented continuous integration workflows to automatically test for the code’s functionality. One of the automatic workflows reviews all the simulations in the examples folder by running them for a few steps and quickly verifying if they are functioning correctly. A second workflow performs regression tests. In these cases, the simulations are performed for the required amount of steps and the simulation outputs are compared against pre-computed reference values. Together, these workflows guarantee that any new modification to the code neither alters the results of simulations nor breaks the provided examples. Additional workflows, less critical for the functioning of the code, focus on checking the formatting syntax using flake8 (https://flake8.pycqa.org/) and black (https://github.com/psf/black), or managing the updates of PyPI releases.
Modifications and additions to i-PI:
To add or modify the source code, contributors must create a pull request in the GitHub repository of i-PI  https://github.com/i-pi/i-pi. The contributors must ensure that all the automatic tests pass successfully. When introducing a new feature, contributors are expected to provide a minimal example that runs with built-in drivers, along with the corresponding regression test to guarantee the long-term stability of the new feature. After these criteria are met, the actual code review starts. During the review process, the code managers might ask the contributor to justify some design decisions, modify part of the code contribution, improve the documentation, etc. Finally, when all the requested changes are fulfilled, the new addition is merged into the main branch of the code.
II.3 How to implement an i-PI  interface in my code?
For softwares that provide a Python interface to directly compute energy and forces for a given structure, such as several machine-learning potentials, it is often possible to create an interface with a few dozen lines of code using the provided Python driver, available in the drivers/py/ folder. In the rest of the cases, the simplest method to implement the communication interface is to identify in the target client software an existing loop that updates nuclear positions as it is done in any geometry optimization or molecular dynamics algorithm. The beginning and the end of this loop are the places where the client code will receive the atomic coordinates from i-PI  and send back energy, forces, stress tensor and other requested quantities (see section IV.2), respectively. In several scenarios, such as the one dealing with electronic structure codes or some MLIPs, the client code would have to carry out the appropriate initialization routines before the client establishes a connection with the i-PI server. Given that such initialization requires knowledge of the types of atoms (that are not communicated by i-PI), it is often necessary to provide a separate description of the structure in the format expected by the client code. Most of the electronic structure clients offer the possibility of using the converged density matrix/wavefunction from a previous step as the initial guess for the subsequent calculation. Consequently, the self-consistent cycle converges as quickly using i-PI as it would with the same algorithm implemented within the client code. Examples of interface implementations in Python and Fortran can be found in the driver folder provided with the code, while a C++ implementation can be found in the LAMMPS code.[37]
II.4 Main Features
We conclude this section by listing the main features available in the 3.0 release:

### v2

# i-PI 3.0: a flexible and efficient framework for advanced atomistic simulations

@@ Abstract
Atomic-scale simulations have progressed tremendously over the past decade, largely due to the availability of machine-learning interatomic potentials. These potentials combine the accuracy of electronic structure calculations with the ability to reach extensive length and time scales. The i-PI package facilitates integrating the latest developments in this field with advanced modeling techniques, thanks to a modular software architecture based on inter-process communication through a socket interface. The choice of Python for implementation facilitates rapid prototyping but can add computational overhead. In this new release, we carefully benchmarked and optimized i-PI for several common simulation scenarios, making such overhead negligible when i-PI is used to model systems up to tens of thousands of atoms using widely adopted machine learning interatomic potentials, such as Behler-Parinello, DeePMD and MACE neural networks. We also present the implementation of several new features, including an efficient algorithm to model bosonic and fermionic exchange, a framework for uncertainty quantification to be used in conjunction with machine-learning potentials, a communication infrastructure that allows deeper integration with electronic-driven simulations, and an approach to simulate coupled photon-nuclear dynamics in optical or plasmonic cavities.
@@ I Introduction
The atomic-scale modeling of molecules and materials is a constantly evolving field driven by the interest of researchers in discovering new fundamental phenomena, interpreting complex experimental results, and, together with an increasing number of commercial enterprises, improving, discovering, and developing new materials.[1, 2, 3, 4, 5, 6] Several computational developments in recent years extended dramatically the problems that can be addressed by atomistic simulations, far exceeding the steady improvement of the underlying hardware. Arguably, the largest contribution to such fast progress can be attributed to the development of machine learning interatomic potentials (MLIPs). MLIPs approach the accuracy of the underlying ab initio data, but at a small fraction of the computational cost. They have allowed access to system sizes and time scales which were previously only accessible to empirical interatomic potentials.[7, 8, 9, 10] This has driven a renewed interest in complex simulation algorithms that were thought inapplicable for studies that require a first-principles description of interatomic interactions, such as the investigation of catalytic reactions [11, 12, 13] or of the thermodynamic stability of materials.[14, 15, 16, 17] However, the lack of easy-to-use implementations of such approaches has become an important factor that hinders a more widespread adoption
@@
Since its first release in 2014 [18], one of i-PI’s primary objectives has been “lowering the implementation barrier to bring state-of-the-art sampling and atomistic modeling techniques to ab initio and empirical interatomic potentials programs”. The first version of i-PI was designed for path integral molecular dynamics (PIMD) simulations, hence the code name. Since then, and thanks to a community effort, it has evolved into a “universal force engine for advanced molecular simulations”, including both standard and advanced simulations algorithms, as described in the following sections.[19]
@@
In the last few years, our efforts have adapted to the changing needs of the field. While continuing with the implementation of new and advanced features, we have improved the scalability of the code backend, as required by newly available and computationally efficient MLIPs. We also adopted automatic testing workflows to strengthen code stability and enhance our reproducibility standards.
@@
i-PI’s capabilities as a versatile tool for atomistic simulations can be demonstrated by the wide range of applications in which it has been employed. These examples range from the dynamics of charge density wave materials [20], interpretation of neutron scatter experiments,[21] and nuclear motion associated to electronic excitations,[22, 23, 24] to the elucidation of the thermodynamic properties of aqueous electrons,[25, 26, 27] the calculation of first principles phase diagrams,[28, 29, 30] and the study of aqueous solutions and interfaces.[31, 32, 33, 34, 35, 36] Moreover, by virtue of its modular structure, i-PI is already being used as a simulation engine with newly developed MLIP architectures [37, 9, 38] making it easy to use them in combination with advanced sampling techniques. It can also be used as a toolbox to combine different simulation “modes” to design sophisticated modeling workflows that involve multiple types of interatomic potentials, free-energy-sampling software[39] as well as post-processing of trajectory data.[40]
@@
This manuscript reports on the progress made in the last 5 years since the release of version 2.0, by describing the newly adopted maintenance strategies, the improved performance, and the newly-implemented features. The manuscript is organised as follows: In Sec. II, we start with a short review of the code architecture, followed by a description of the improvements made to the code’s robustness and usability and a list of the main features available in this release. In Sec. III, we present a detailed discussion of the code efficiency for production and large-scale simulations while in Sec. IV, we describe some selected new features since the last code release discussing their implementation and examples of their application. In Sec. V, we showcase how the modularity and flexibility of i-PI help the implementation of advanced simulation setups that would otherwise require new codes to be written and Sec. VI concludes the article.
@@ II Program Overview
i-PI  is a force engine that operates with a client-server paradigm. i-PI  acts as a server in command of the evolution of the nuclear positions while one or more instances of the client code handle the evaluation of energy, forces, and possibly other properties for individual configurations. Designed to be universal, i-PI’s architecture is agnostic to the identity of the force providers, with a general and flexible backend that accommodates a wide range of client codes.
@@
The code is distributed under a dual GPL and MIT license, and it is written in Python 3, a high-level, general-purpose interpreted programming language known for its emphasis on code readability. This choice facilitates rapid prototyping of new ideas and relatively easy code maintenance when compared to compiled languages traditionally used in scientific computing. Additionally, the availability of highly optimised numerical libraries allows Python-based software to deliver competitive performance. Indeed, even computing-intensive electronic structure codes have released Python implementations in recent years.[41, 42] A detailed discussion of the code efficiency and improvements since the previous release is presented in Sec. III.
@@
The communication between i-PI  and the client codes is carried out through sockets, enabling local connections by shared-memory inter-process communication (UNIX domain sockets) or between two hosts connected to a local network or to the Internet (TCP/IP sockets). During the initialization, i-PI creates the sockets and awaits the connection of one or more clients. When the connection is established, the simulation loop starts: i) i-PI communicates the nuclear positions and the cell lattice vectors to the active clients, ii) the clients evaluate the required properties – typically energies, forces, and stresses – and send these back, and iii) i-PI uses this data to displace the atoms and to calculate the new nuclear positions, restarting the loop. As i-PI  is designed to be client-agnostic, clients must be initialized independently before they can connect to the sockets. Importantly, the identity and the number of atoms must be defined consistently in the client code according to its own input mechanism and remain unchanged throughout the simulation.
@@
A schematic representation of the code structure and the server-client communication is presented in Fig. 1. i-PI is structured in a modular way that represents the underlying physics of the target simulation as faithfully as possible. To achieve this, the code is organized around the System class which encodes all the information related to the physical system, such as the number and identity of atoms, the ensemble to be sampled, the number of replicas in the path integral simulations, the initialization procedure, and the algorithm for evolving the nuclear positions. Forces are managed through the Forces class, which can combine multiple force components, each of which is computed following the strategy specified in a Forcefield class. This two-layer approach is particularly advantageous for algorithms where the total forces and energies are obtained as the sum of terms computed at different accuracy levels,[43, 44, 24, 45, 46, 47] or for different portions of the system.[48] Simulation techniques that require the evolution of many systems simultaneously make use of the SMotion (Systems Motion) class. This class enables the definition of evolution steps that combine the (possibly many) different systems, facilitating, for example, exchanges between system replicas as done in replica exchange simulations, or the application of biases, like in metadynamics. Finally, estimators can be defined within the code or computed by the provided post-processing tools or user-generated scripts. An updated list of code features is reported in Section II.3.
II.1 Code philosophy and development model
Many available client codes:
@@ II.1 Code philosophy and development model
i-PI  is connected to a multitude of electronic structure and machine learning codes, either directly or through its ASE [49] or LAMMPS [50] interfaces. This wide-ranging connectivity means that a single implementation of an advanced simulation method is readily available to all the users of these codes, which significantly amplifies the potential impact of each contribution to i-PI. Furthermore, this capability enables a direct comparison between different client codes within an identical simulation setup.
Ease of implementation of new interfaces:
Adding a new interface to i-PI is simple. For codes that provide a Python interface, this can be implemented by adding a small new file to the Python driver directory (drivers/py/pes/). For compiled software, the driver loop can usually be rather self-contained, modelled after the molecular dynamics or geometry optimization loop. The FORTRAN driver distributed with i-PI provides a good example, as well as utilities to open and communicate over sockets that can be reused in third-party codes, while the LAMMPS[50] implementation of fix_ipi can be used as a template of a C++ interface. In all cases, the structure (atom numbers and types) should be initialized before communication is established, and that i-PI does not guarantee that successive calls provide similar structures (as in a MD trajectory) which might require some care when the force evaluator makes a continuity assumption for neighbor list updates or wavefunction extrapolation.
@@
In scenarios where a developer aims to use a newly implemented algorithm for nuclear motion with different types of codes, such as a density-functional theory (DFT) plane wave code to study periodic systems, or a quantum chemistry code to study gas phase molecules at a higher electronic structure level, two separate implementations would be required. By integrating the algorithm into i-PI, only one implementation is needed and there is only one code to maintain.
@@
The infrastructure of i-PI automatically manages updates of energy, forces, stresses, and other structure-dependent properties by establishing dependencies between physically related variables. This architecture enables developers to concentrate solely on the algorithm’s implementation making the development process more focused and efficient.
Participative development model:
The modular structure of the code simplifies the contribution of self-contained implementations of new methods. These methods are explicitly listed in a dedicated section of the documentation, giving developers visibility and recognition for their contributions. Many of the applications presented in this paper are examples of this model.
Code Maintenance and Long-term Sustainability:
As with most open-source projects, i-PI is developed and maintained by a community of volunteer contributors. Several of the authors of this manuscript, who are core developers and hold permanent or long-term academic positions, are involved in the governance of the code. These people have administrator rights on the main repository, which reduces the risk that the code base will be abandoned. We encourage power users to also help maintain the core infrastructure and the documentation; in preparation for this release, we have dedicated a special focus on improving the development infrastructure, as we discuss in the next Section.
@@ II.2 Code robustness and usability
The increased number of i-PI  users and contributors since the last release required enhancing the maintainability, robustness, and standardization of the code. These improvements involve five distinct aspects:
@@
We refactored the documentation of i-PI, which is now generated using Sphinx, is deployed automatically when new features are merged to the main branch, and is linked to the main code webpage 111i-PI official documentation: https://ipi-code.org/i-pi/index.html. The documentation has been updated to reflect the current status of the code, and a section about how to contribute new features to the code has been added to reduce the barrier for new developers. An automatic parsing of the Python docstrings throughout i-PI ensures a complete listing of all possible keywords and the corresponding explanation, which is automatically updated as soon as new features are implemented and merged in the main branch.
@@
We have completely restructured the provided examples to make them easier to navigate and as intuitive as possible for newcomers. The examples are now organized into three directories, namely features, clients, and hpc_scripts. The features directory displays extensive examples of the different implemented functionalities, all of which can be run locally using the built-in Python or Fortran driver, without requiring additional software installation. The clients directory includes examples tailored to specific driver codes, highlighting the unique syntax and tags needed for an adequate setup when used in connection to i-PI. At the moment, we provide examples for ASE,[49] CASTEP,[52] CP2K,[53] LAMMPS,[50] DFTB+,[54] ffsGDML,[55] FHI-aims,[56] librascal,[57] Quantum ESPRESSO,[58, 59, 60] VASP,[61] elphmod,[62] and YAFF.[63] We reiterate that many other codes can also be connected to i-PI through their ASE and LAMMPS interfaces. Finally, motivated by the number of user queries on this topic, the hpc_scripts directory provides examples with the corresponding submission scripts to run the simulations on high-performance computing platforms.
@@
We have added a demos directory that contains examples of more complicated setups, to illustrate the application of i-PI  in realistic simulation scenarios, combining several of the features showcased in the examples directory. Some of these demonstration examples also cover some of the use of the pre- and post-processing scripts accompanying the code. Users specifically interested in learning the theory of path integral simulations and obtaining hands-on practice using i-PI  are referred to the online courses developed with this goal. [^2] The “Path Integral Methods in Atomistic Modelling” is hosted at https://app.courseware.epfl.ch/learning/course/course-v1:EPFL+X+2022/home and code tutorials at https://github.com/i-pi/piqm2023-tutorial/
@@
We have implemented continuous integration workflows to automatically test for the code’s functionality. One automatic workflow reviews all the simulations in the examples directory by running them for a few steps and quickly verifying if they are functioning. A second workflow performs regression tests. In these cases, the simulations are performed for the required amount of steps and the simulation outputs are compared against pre-computed reference values. Together, these workflows guarantee that any new modification to the code neither alters the results of simulations that are tested nor breaks the provided examples. Additional workflows, less critical for the functioning of the code, focus on checking the formatting syntax using Flake8 333 See https://flake8.pycqa.org/ for more information about Flake8 and Black 444 See https://github.com/psf/black for more information about Black , or managing the updates of PyPI releases.
Developers’ documentation and workflows
We have included and updated the development guidelines, that are discussed in the main documentation, as well as in the CONTRIBUTING.md file in the main repository. Contributors are encouraged to use the standard git workflows, such as creating a fork and opening a pull request, and expect code review that will involve both the correctness of the algorithm and the consistency with i-PI’s coding style and quality.
II.3 Main Features
We conclude this Section by listing the main features currently available in i-PI. A comprehensive list of features, indicating the main contributors, is maintained as part of the documentation.
II.3.1 Features available in the 2.0 release: